OpenStudy (anonymous):

need help to simplify the following -15 xsquared yto the third --------------------------- 12 xto the third ysquared

OpenStudy (amistre64):

what should your first step be?

OpenStudy (anonymous):

\[\frac{-15x^2y^3}{12x^3y^2}=\frac{-5y}{4x}\]

OpenStudy (amistre64):

-15 x^2 y^3 ------------- 12 x^3 y^2 what do you know about these "terms"?

OpenStudy (anonymous):

@ nadeem what 're Ur reason to think like that?

OpenStudy (amistre64):

when we multiply like bases; the exponents "add" together: 5^2 5^8 = 5^(2+8) = 5^10 when we do the opposite of multiplication (which is division) we do the opposite with the exponents as well; which is subtraction: 5^8 ---- = 5^(8-2) = 5^6 5^2 Do you ubderstand that "rule" of bases and exponents?

OpenStudy (anonymous):

my new math knowledge amistre,,,thx...

OpenStudy (amistre64):

yw :)

OpenStudy (anonymous):

you can break it up and use exponential rules: \[(\frac{-15}{12})(\frac{x^2}{x^3})(\frac{y^3}{y^2}) \rightarrow (\frac{-5}{4})(\frac{1}{x})(\frac{y}{1})\rightarrow \frac{-5y}{4x}\]

OpenStudy (anonymous):

yeah i see thx..